# NetMamba: Efficient Network Traffic Classification via Pre-training Unidirectional Mamba

### v1

@@ VII-D Ablation Study
To further validate the design of NetMamba, we conducted an ablation study to assess the contribution of each component across six public datasets. The corresponding results are presented in Table VII.
@@ VII-D1 Model-level Ablation
Initially, we replaced all unidirectional Mamba blocks in NetMamba with bidirectional ones. The experimental results revealed a slight performance decline across all datasets except for ISCXVPN2016. This suggests that unidirectional Mamba is well-suited for processing network traffic data, given that packets are transmitted sequentially and earlier packets possess limited information about subsequent ones. Moreover, incorporating bidirectional or even omnidirectional Mamba blocks introduces additional computational and memory overheads due to extra scan passes, ultimately reducing efficiency. Thus, unidirectional Mamba stands out as the preferable choice.
While positional information is inherently preserved in sequence models such as Mamba, eliminating explicit positional embedding still results in a reduction in accuracy ranging from 0.03% to 1.35% across all datasets except for ISCXTor2016. Therefore, incorporating positional information enhances the model’s capability for better traffic comprehension.
@@ VII-D2 Data-level Ablation
When omitting all header bytes within a packet, there is a significant decrease in classification performance, with accuracy dropping between 14.51% and 48.54%. This underscores the importance of key fields within headers of the network layer, transport layer, or other upper layers.
Regarding packet payloads, the ablation results demonstrate accuracy drops ranging from 0.26% to 8.29% across five datasets, with a slight increase of 0.07% on the ISCXTor2016 dataset. This underscores the contribution of potential plaintext and specific encrypted payloads to improved traffic understanding.
@@ VII-E Few-Shot Evaluation
To validate the robustness and generalization abilities of NetMamba, we conduct few-shot evaluations on four datasets, with labeled data size set to 10%, 40%, 70%, and 100% of the full training set (comprising 80% of the total data). As shown in Figure 5, the three pre-trained models, NetMamba, YaTC, and ET-BERT generally outperform other supervised methods under few-shot settings. While conventional machine learning methods like FlowPrint and AppScanner show some robustness to limited labeled data, their classification performance varies significantly across different datasets. Although the supervised TFE-GNN model performs comparably to the pre-trained models with the full training dataset, its performance drops considerably with smaller training data sizes. Thus, pre-trained models demonstrate superior robustness and generalization capabilities due to their ability to extract high-quality traffic representations from large amounts of unlabeled data, thereby reducing the dependence on labeled data.
@@ VIII Conclusion and Future Work
In this paper, we introduce NetMamba, a novel pre-trained state space model designed for efficient network traffic classification. To enhance model efficiency while maintaining performance, we utilize the unidirectional Mamba architecture for traffic sequence modeling and develop a comprehensive representation scheme for traffic data. Evaluation experiments on six public datasets demonstrate the superior effectiveness, efficiency, and robustness of NetMamba.
Beyond classical traffic classification tasks, the comprehensive representation scheme and refined model design enable NetMamba to address broader tasks within the network domain, such as quality of service prediction and network performance prediction. However, the current implementation of NetMamba depends on specialized GPU hardware, which limits its deployment on real-world network devices. In the future, we plan to explore solutions to implement NetMamba on resource-constrained devices.

### v3

@@ VII-D Ablation Study
To further validate both the model design and the traffic representation scheme of NetMamba, we conducted ablation studies to assess the contribution of each component across six public datasets. The results are presented in Table VI.
@@ VII-D1 Model-level Ablation
Initially, we replaced all unidirectional Mamba blocks in NetMamba with bidirectional ones used by [12]. The experimental results revealed a slight performance decline across all datasets except for ISCXVPN2016. This suggests that unidirectional Mamba is well-suited for processing network traffic data, given that packets are transmitted sequentially and earlier packets possess limited information about subsequent ones. Moreover, incorporating bidirectional or even omnidirectional Mamba blocks introduces additional computational and memory overheads due to extra scan passes, ultimately reducing efficiency. Thus, unidirectional Mamba stands out as the preferable choice.
Following [30], we substituted the original unidirectional Mamba block with a cascading structure where each inner block processes data of different granularity. We observed a notable drop in classification performance across all datasets, indicating that this complex structure is inferior to processing sequential traffic data.
To clarify the source of NetMamba’s classification performance gains over existing Transformer-based models, we further evaluated two Transformer-based ablation variants, NT-Vanilla and NT-Linear. The results show that NT-Vanilla performs similarly to NetMamba across all datasets, suggesting that our proposed traffic representation scheme plays a key role in enhancing classification performance. With this representation scheme, a linear-time Mamba-based classifier achieves classification performance comparable to that of a quadratic-time Transformer-based model. However, NT-Linear, due to information loss from the overcompression of attention mechanisms, exhibits unstable performance and falls significantly behind NetMamba on three datasets.
While positional information is inherently preserved in sequence models such as Mamba, eliminating explicit positional embedding still results in a reduction in accuracy ranging from 0.03% to 1.35% across all datasets except for ISCXTor2016. This suggests that reinforced positional information aids the model in capturing correlations within sequential traffic data.
@@ VII-D2 Data-level Ablation
When all header bytes in a packet are omitted, classification performance declines significantly, with accuracy dropping by 14.51% to 48.54%. This highlights the critical role of key fields within packet headers—such as port number, protocol, and packet length—which have been proven effective in traffic classification [4, 38, 39].
Regarding packet payloads, the ablation results show accuracy drops ranging from 0.26% to 8.29% across five datasets. This underscores the contribution of potential plaintext and specific encrypted payloads to improved traffic understanding.
@@ VII-E Few-Shot Evaluation
To validate the robustness and generalization abilities of NetMamba, we conduct few-shot evaluations on four datasets, with labeled data size set to 10%, 40%, 70%, and 100% of the full training set (comprising 80% of the total data). Specifically, we adopt a leave-one-out approach on the pre-training datasets to assess the transfer learning capability of the pre-trained models. In detail, the dataset used for fine-tuning is excluded from the pre-training datasets. As shown in Figure 5, the three pre-trained models, NetMamba, YaTC, and ET-BERT generally outperform other supervised methods under few-shot and leave-one-out settings. While conventional machine learning methods like FlowPrint and AppScanner show some robustness to limited labeled data, their classification performance varies significantly across different datasets. Although the supervised TFE-GNN model performs comparably to the pre-trained models with the full training dataset, its performance drops considerably with smaller training data sizes. Thus, pre-trained models demonstrate superior robustness and generalization capabilities due to their ability to extract high-quality traffic representations from large amounts of unlabeled data, thereby reducing the dependence on labeled data.
@@ VIII Conclusion and Future Work
In this paper, we introduce NetMamba, a novel pre-trained state space model designed for efficient network traffic classification. To enhance model efficiency while maintaining performance, we utilize the unidirectional Mamba architecture for traffic sequence modeling and develop a comprehensive representation scheme for traffic data. Evaluation experiments on six public datasets demonstrate the superior effectiveness, efficiency, and robustness of NetMamba. Beyond classical traffic classification tasks, the comprehensive representation scheme and refined model design enable NetMamba to address broader tasks within the network domain, such as quality of service prediction and network performance prediction. However, the current implementation of NetMamba depends on specialized GPU hardware, which limits its deployment on real-world network devices. In the future, we plan to explore solutions to implement NetMamba on resource-constrained devices.